OpenStudy (anonymous):

Solve the exponential equation using logarithms, rounding the answer to the nearest thousandth (3 decimal places). Show your work. 2 = 7^x+3

OpenStudy (anonymous):

i'm assuming you mean 2 = 7^(x+3) ?? or is it 2 = 7^x + 3

OpenStudy (anonymous):

no its 2=7^x+3

OpenStudy (anonymous):

no solution then... boy that was easy huh?

OpenStudy (anonymous):

ok.. since its the second one i wrote down, if you subtract 3 from both sides, you'll get: -1 = 7^x correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now, 7^x will never ever forever be negative, no matter what x value you put in...

OpenStudy (anonymous):

in other words, 7^x will always be a positive number. so there is no solution to your original equation.

OpenStudy (anonymous):

ok sounds good

OpenStudy (anonymous):

ok... have a good one man.. :)